OpenStudy (anonymous):

a parent died and left an estate to four children. one inherited 1/5, the second inherited 3/10, the third inherited 2/5. how much did the fourth child inherit in the form of a fraction?

OpenStudy (anonymous):

3/10.

OpenStudy (anonymous):

sorry i mean 1/10

OpenStudy (hoblos):

=1- 1/5 - 3/10 - 2/5 =10/10 - 2/10 - 3/10 - 4/10 = 1/10

OpenStudy (anonymous):

thank you for showing me the process hoblos

OpenStudy (anonymous):

i got 1/10 also